OpenStudy (anonymous):

Evaluate the dot product of (5,7) and (-8,2).

OpenStudy (anonymous):

the dot product of two vectors will give you back a number (scalar, not another vector). so (5,7) dot (-8,2) = 5*7 + -8*2 = 35 -16 = 19